OpenStudy (anonymous):

Hi, I need help with the steps! Use a definite integral to find an expression that represents the area of the region between the given curve and the x-axis on the interval [0,b]. y=6pi x^2

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

x is restricted to 0 and b, then we sub this domain to find region of y. y then must be between 0 and 6pib^2.

OpenStudy (anonymous):

area is just \[\int\limits\limits_{0}^{b} \int\limits_{0}^{6\pi b^2} 1 dydx\]

OpenStudy (anonymous):

\[\int\limits_{0}^{b}y dx =\int\limits_{0}^{b}6\pi x^2 dx=6\pi \int\limits_{0}^{b} x^2 dx\] \[=6\pi [\frac{x^3}{3}]_{0}^{b} =\frac {6\pi}{3}[{x^3}]_{0}^{b}= 2\pi [b^3-0^3] = 2\pi b^3\]

OpenStudy (anonymous):

soz, the upper bound of y region should be 6 pix^2.

OpenStudy (anonymous):

SO the final answer is 2pi b^3 @dpasingh